OpenStudy (he66666):

Grade 12 math: Trig identities/equations? the depth of the ocean at a swim buoy can be modelled by the function d(t) = 4 + 2sin[(pi/6)t], where d is the depth of the water in metres and t is the time in hours, if t is between 0 and 24. Consider a day when t=0 represents midnight. Determine when the depth of water is 3m. answer: t=7, 11, 19, and 23 How do you solve this question? I don't get it.. I got until sin[(pi/6)t] = -1/2...

OpenStudy (anonymous):

d = 3 = 4 +sin(pi*t/6) sin (pi*t/6) = -1 pi*t/6 = 3*pi/2 or 7*pi/2 t = 9 or t = 21

OpenStudy (he66666):

There's a 2 in front of sin(pi*t/6) though :S

OpenStudy (anonymous):

Dang: pi*t/6 = 11pi/6 or 7pi/6 or 23pi/6 or 19pi/6 t = 7,11,19, or 23

OpenStudy (he66666):

Thanks :)